OpenStudy (anonymous):

Please help An apartment building contains 12 units consisting of one- amd two bedroom apartments that rent for $360 and $450 per month respectively. When All units are rented, the total monthly rental is $4,950. what is the number of two bedroom apartments? A.3 B.4 C. 5 D. 6 E. 7

OpenStudy (gabebae):

x+y=12 y=12-x 360x+450y=4950 360x+450(12-x)=4950 360x+5400-450x=4950 -90x=-450 x=5 Answer is C.5

OpenStudy (anonymous):

there are 5 one bedroom apts and 7 two bedroom apts

OpenStudy (anonymous):

so the answer is E

OpenStudy (gabebae):

Which answer is correct? I'd like to know

OpenStudy (anonymous):

I thought you was sure You did all the equation xD

OpenStudy (anonymous):

the equations are correct. it's just that @GabeBae solved for X and X = to the 1 room apts you need to solve for Y which is equal to the 2 room apts

OpenStudy (anonymous):

That is why it is 7 and not 5 ;)

OpenStudy (anonymous):

Which is the x number you have to solve for?

OpenStudy (anonymous):

"what is the number of two bedroom apartments?"

OpenStudy (anonymous):

the two bedroom apartments are Y not X.

OpenStudy (anonymous):

2

OpenStudy (anonymous):

x + y = 12 360x + 450y = 4,950 x = 12 - y 360(12 - y) + 450y = 4950 -360y + 4320 + 450y = 4950 90y = 630 y = 7

OpenStudy (anonymous):

since 2 bedroom apartments will cost more than a one bedroom we can see that y is the variable you want to solve for

OpenStudy (anonymous):

ok thanks :)